The demand equation for the Drake GPS Navigator is
x + 4p − 728 = 0,
where x is the quantity demanded per week and p is the wholesale unit price in dollars. The supply equation
x − 20p + 1000 = 0,
where x is the quantity the supplier will make available in the market each week when the wholesale price is p dollars each. Find the equilibrium quantity and the equilibrium price for the GPS Navigators.
equilibrium quantity units
equilibrium price $

Answer by ikleyn(52777)

.
The demand equation for the Drake GPS Navigator is
x + 4p − 728 = 0,
where x is the quantity demanded per week and p is the wholesale unit price in dollars.

The supply equation is
x − 20p + 1000 = 0,
where x is the quantity the supplier will make available in the market each week
when the wholesale price is p dollars each.
Find the equilibrium quantity and the equilibrium price for the GPS Navigators.
~~~~~~~~~~~~~~~~~~

They want you find the values x and p that satisfy the two given equations SIMULTANEOUSLY.

    x + 4p − 728 =   0,     (1)
    x - 20p + 1000 = 0.     (2)


It is easy.  From equation (1), subtract equation (2).
The terms with "x" wil cancel each other, and you will get

       4p - (-20p) - 728 - 1000 = 0,

or

      24p = 1728,  --->  p = 1728/24 = 72.


Next, from equation (1) you have, substituting p= 72 there

    x + 4*72 - 728 = 0

    x - 440 = 0

    x = 440.


ANSWER.  The solution is  x= 440,  p= 72.

Solved, with explanations.